Ghana brewery limited produces two brands of products, product A and B. Their demand functions are, i. Q1=14-0.25p1 ii.Q2=24-0.5p2 respectively. Their joint cost function = TC=Q1^2 +5Q1.Q2^2+Q2^2. Estimate price and output level that will maximize profit for the company? (Where ^ represents exponent. For example Q1^2, meaning Q1 exponent 2.)
1
Expert's answer
2017-12-25T08:14:07-0500
Q1 = 14 – 0.25*P1
Q2 = 24 – 0.5*P2
Total Revenue = P1*Q1 + P2*Q2
= (56 – 4*Q1)Q1 + (48 – 2*Q2)Q2
= 56*Q1 + 48*Q2 – 4*Q12 – 2*Q22
Total Cost = Q12 + 5*Q1*Q22 + Q22
P (Profit Function) = TR – TC
P = 56*Q1 + 48*Q2 – 5*Q12 – 3*Q22 – 5*Q1*Q22
Differentiating with respect to Q1 and Q2 and equating it to 0.
56 – 10*Q1 – 5*Q22 = 0
48 – 6*Q2 – 10*Q1*Q2 = 0
Substituting 10*Q1 from first equation to the second to find optimum Q2.
-10*Q1 = 5*Q22 – 56
So, 48 – 6*Q2 + 5*Q23 – 56*Q2 = 0
48 – 62*Q2 + 5*Q23 = 0
This equation does not have integral solutions but it has a solution very close to 3. So, we will take that as an optimum value of Q2.
So, Q2 = 3
10*Q1 = 11 which is close to 1 if we took 3.05 instead of 3 for Q2. So, optimum Q1 = 1.
P1 = 52 and P2 = 42.